# Tailored Emotional LLM-Supporter: Enhancing Cultural Sensitivity

### v1

@@ E.2 Additional Crowd Evaluation Results
We present additional crowd evaluation results in Table 15. Our results show that explicitly providing cultural signals ( / ) and using a combined strategy () are the most preferred strategies. Further, humans and models agree majority of the time for the best adaptation strategies (shaded cells in Table 15), even though Kendall’s might be low.
| Model | Arabic | Chinese | German | Jewish | ||||
| Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | Emo. | Cult. | |
| Aya-Expanse-8B | +cga | +cga | +a | +a | +a | +a | +a | +a |
| Qwen-2.5-7B | +g | +cga | +a | +cga | +cga | +cga | +a | +cga |
| Llama-3.1-8B | +cga | +cga | +g | +cga | +g | +a | +cga | +cga |
@@ Appendix F Examples
We show annotated examples from CultureCare in Tables 17 and 18. These examples show culturally nuanced mental health posts, the corresponding human, and (adapted) LLM-generated responses. Each post is annotated with colour-coded signals for distress messages and their intensity, cultural signals and their categories, and support strategies.
@@ Appendix G Details about CCCI-R
CCCI-R LaFromboise et al. (1991a, b) is a commonly used and well-studied 20-item questionnaire developed to assess the cross-cultural counselling competency of professional counsellors and counsellors in training. In this work, we adapted a subset of items from the CCCI-R that align with the goals of this paper, as many items in the original questionnaire focus on counsellors’ professional responsibility and in-person non-verbal communication.
@@ H.2 Inter-annotator Agreement
We show the inter-annotator agreement results in Table H.2. The original (Orig.) numbers show the agreement on our original rating scales (as shown in Figure 22). The annotators have a high agreement on what is an overly positive or an overly negative response. For the rest of the criteria, we see moderate to low agreement. After closer inspection, we notice that the clinicians, in many cases, align in attitude towards a certain response, but do not align in the level of agreement. That is, they might both think that a certain response shows some emotional empathy, but one would choose to “Strongly Agree” and the other to just “Agree”. We verify this observation by reporting grouped (Grou.) agreement numbers. We find that collapsing categories to allow for partial agreement improves the agreement levels of all metrics.
@@ H.3 Results
Our numerical results from the expert evaluation are in Figure 23. Overall, LLMs consistently avoid extreme (both negative and positive) emotional tones when providing responses. The empathy and emotional regulation metrics further highlight LLM strengths. For cognitive empathy, the majority (78%) of LLM responses were rated as “Strongly agree” (30%) or “Agree” (48%). Emotional empathy showed similar results, 34% ”Strongly agree” and 46% “Agree”. Emotion regulation followed the same pattern, 36% “Strongly agree” and 44% “Agree”. Human ratings in these categories were more distributed, with lower agreement levels and higher disagreement. On cultural understanding, 67% of LLM responses were rated “yes”, 31% “partially”, and only 2% “no”. When evaluating LLM usefulness for training new psychologists, 46% were seen as strong examples, 42% as promising with needing revision, and only 12% as not suitable, suggesting LLMs offer reliable, culturally aware, and emotionally attuned outputs with clear training potential.

### v2

@@ Appendix D Additional Automatic Evaluation Results
In this paper, we mainly focused on examining models in the 7-8B parameter range, given their popularity, performance, and suitability for agentic systems Belcak et al. (2025), while still remaining competitive. Here, we provide additional results for two larger models—Llama-3.1-70B and Qwen-2.5-72B—in Table 16. Here, we observe similar patterns as with small models. Overall, the +guided strategy is more effective at providing winning emotional support, while +cga is better attuned to cultural sensitivity. The overall winning strategy with Qwen-2.5-72B is +cga, whereas for Llama-3.1-70B it is +guided. Simple culture-informed role-playing (+culture) does not substantially improve cultural awareness compared to +cga. Interestingly, the best overall average automatic evaluation scores for both Llama-3.1-70B (4.05) and Qwen-2.5-72B (4.12) are lower than that of Aya-Expanse-8B (4.51). This indicates that a larger model may not perform better for our specific task, necessitating further research.
| Model | Arabic | Chinese | German | Jewish | Average | ||||||
| Llama-3.1-70B | |||||||||||
| redditor | 3.82 | 3.47 | 4.07 | 3.53 | 3.86 | 3.11 | 4.39 | 4.07 | 4.03 | 3.54 | 3.79 |
| +culture | 3.75 | 3.97 | 4.01 | 4.13 | 3.94 | 3.14 | 4.39 | 4.18 | 4.02 | 3.86 | 3.94 |
| +guided | 4.25 | 3.99 | 4.24 | 4.01 | 4.05 | 3.13 | 4.47 | 4.27 | 4.25 | 3.85 | 4.05 |
| +annotation | 4.08 | 3.92 | 4.05 | 3.94 | 3.78 | 3.12 | 4.48 | 4.14 | 4.10 | 3.78 | 3.94 |
| +cga | 3.74 | 4.23 | 4.24 | 4.32 | 3.69 | 3.20 | 4.39 | 4.35 | 4.01 | 4.02 | 4.02 |
| Qwen-2.5-72B | |||||||||||
| redditor | 4.27 | 3.30 | 4.22 | 3.23 | 4.14 | 3.04 | 4.33 | 3.86 | 4.24 | 3.36 | 3.80 |
| +culture | 4.05 | 3.58 | 4.20 | 3.73 | 4.33 | 3.21 | 4.64 | 4.02 | 4.31 | 3.63 | 3.97 |
| +guided | 4.49 | 3.58 | 4.60 | 3.58 | 4.71 | 3.13 | 4.76 | 3.82 | 4.64 | 3.53 | 4.08 |
| +annotation | 4.18 | 3.38 | 4.34 | 3.47 | 4.17 | 3.02 | 4.48 | 3.80 | 4.29 | 3.41 | 3.85 |
| +cga | 4.37 | 3.93 | 4.28 | 3.93 | 4.37 | 3.33 | 4.62 | 4.14 | 4.41 | 3.83 | 4.12 |
@@ E.2 Additional Crowd Evaluation Results
We present additional crowd evaluation results in Table 5. Our results show that explicitly providing cultural signals ( / ) and using a combined strategy () are the most preferred strategies. Further, humans and models agree on the majority of the time for the best adaptation strategies (shaded cells in Table 5), even though Kendall’s might be low.
@@ Appendix F Examples
We show annotated examples from CultureCare in Tables 18 and 19. These examples show culturally nuanced mental health posts, the corresponding human, and (adapted) LLM-generated responses. Each post is annotated with colour-coded signals for distress messages and their intensity, cultural signals and their categories, and support strategies.
@@ Appendix G Details about CCCI-R
CCCI-R LaFromboise et al. (1991b) is a commonly used and well-studied 20-item questionnaire developed to assess the cross-cultural counselling competency of professional counsellors and counsellors in training. In this work, we adapted a subset of items from the CCCI-R that align with the goals of this paper, as many items in the original questionnaire focus on counsellors’ professional responsibility and in-person non-verbal communication.
@@ H.2 Inter-annotator Agreement
We show the inter-annotator agreement results in Table 20. The original (Orig.) numbers show the agreement on our original rating scales (as shown in Figure 22). The annotators have a high agreement on what is an overly positive or an overly negative response. For the rest of the criteria, we see moderate to low agreement. After closer inspection, we notice that the clinicians, in many cases, align in attitude towards a certain response, but do not align in the level of agreement. That is, they might both think that a certain response shows some emotional empathy, but one would choose to “Strongly Agree” and the other to just “Agree”. We verify this observation by reporting grouped (Grou.) agreement numbers. We find that collapsing categories to allow for partial agreement improves the agreement levels of all metrics.
@@ H.3 Results
As per Figure 3, LLMs consistently avoid extreme (both negative and positive) emotional tones when providing responses. The empathy and emotional regulation metrics further highlight LLM strengths. For cognitive empathy, the majority (78%) of LLM responses were rated as “Strongly agree” (30%) or “Agree” (48%). Emotional empathy showed similar results, 34% ”Strongly agree” and 46% “Agree”. Emotion regulation followed the same pattern, 36% “Strongly agree” and 44% “Agree”. Human ratings in these categories were more distributed, with lower agreement levels and higher disagreement. On cultural understanding, 67% of LLM responses were rated “yes”, 31% “partially”, and only 2% “no”. When evaluating LLM usefulness for training new psychologists, 46% were seen as strong examples, 42% as promising with needing revision, and only 12% as not suitable, suggesting LLMs offer reliable, culturally aware, and emotionally attuned outputs with clear training potential.
H.4 Real-world Implementation
In this work, we demonstrated through expert evaluation that LLM-generated responses have strong potential for training psychologists in culturally competent therapy (see Figure 3). Building on these findings, there are several concrete ways to apply them in real-world educational settings. Instructors could use LLM outputs as case studies for critique, asking students to evaluate emotional support quality and cultural sensitivity. Students could compare unadapted versus culturally adapted responses to understand how cultural signals shape communication. LLMs could also simulate culturally diverse patients for role-playing exercises, allowing students to practice responding in real time or revising responses to improve empathy and cultural competence. Additional exercises might include modifying LLM outputs to align with best practices or generating follow-up questions to elicit culturally relevant information. With professional oversight, these structured activities provide hands-on opportunities to translate our findings into practical training for future psychologists. Implementing these applications is beyond the scope of this paper, as it would require close collaboration with psychologists who have access to a large number of students and the necessary training resources.
@@ Appendix I Distributions of Support Message Categories
We qualitatively evaluate responses generated by Aya-Expanse-8B, which was preferred by most cultural groups. To analyze the types of emotional support, we use Llama-3.1-70B to categorize responses based on the framework from Liu et al. (2021). We manually inspected a random sample of the categories assigned by Llama-3.1-70B and found no inconsistencies. Figure 23 shows the distribution of emotional support message categories of Aya-Expanse-8B’s responses across cultures for the strategy. Additional histograms for other adaptation strategies are included in Figure 24.
Overall, model-generated responses use a mix of reflection of feelings, affirmation, suggestions, and information, which is distinctly different from human support online that primarily offers suggestions. This pattern is consistent across cultures and adaptation strategies. Compared to natural human responses, LLM-generated ones tend to be more verbose and structured, particularly when using compound support strategies (). Tables 18 and 19 (Appendix F) contain specific response examples.